### v1

# Reinforced Latent Reasoning for LLM-based Recommendation

Abstract
Large Language Models (LLMs) have demonstrated impressive reasoning capabilities in complex problem-solving tasks, sparking growing interest in their application to preference reasoning in recommendation systems. Existing methods typically rely on fine-tuning with explicit chain-of-thought (CoT) data. However, these methods face significant practical limitations due to (1) the difficulty of obtaining high-quality CoT data in recommendation and (2) the high inference latency caused by generating CoT reasoning. In this work, we explore an alternative approach that shifts from explicit CoT reasoning to compact, information-dense latent reasoning. This approach eliminates the need for explicit CoT generation and improves inference efficiency, as a small set of latent tokens can effectively capture the entire reasoning process. Building on this idea, we propose Reinforced Latent Reasoning for Recommendation (LatentR3), a novel end-to-end training framework that leverages reinforcement learning (RL) to optimize latent reasoning without relying on any CoT data. LatentR3 adopts a two-stage training strategy: first, supervised fine-tuning to initialize the latent reasoning module, followed by pure RL training to encourage exploration through a rule-based reward design. Our RL implementation is based on a modified GRPO algorithm, which reduces computational overhead during training and introduces continuous reward signals for more efficient learning. Extensive experiments demonstrate that LatentR3 enables effective latent reasoning without any direct supervision of the reasoning process, significantly improving performance when integrated with different LLM-based recommendation methods. Our codes are available at https://anonymous.4open.science/r/R3-A278/.
1 Introduction
Enhancing the reasoning capabilities of Large Language Models (LLMs) has been a central research objective since the LLMs’ emergence, with recent advances—such as DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (OpenAI, 2024)—fueling a surge of interest in this direction. By being trained or architected to reason more deliberately through techniques like chain-of-thought (CoT) (Wei et al., 2022) reasoning, LLMs have demonstrated remarkable progress in tackling complex real-world problems (Xu et al., 2025; Shao et al., 2024), rivaling or even surpassing human PhD-level performance in certain cases (OpenAI, 2024). These developments have generated growing enthusiasm for applying LLM reasoning to downstream tasks, including one key AI application for information access—recommender systems (Tsai et al., 2024; Bismay et al., 2024). For recommendation, the core lies in reasoning user preferences from historical behaviors (Tsai et al., 2024), which only implicitly indicate preference. This is well aligned with the strengths of LLM reasoning, offering the potential to unlock new paradigms for personalized recommendation.
Existing approaches to LLM reasoning for recommendation typically rely on explicit textual chain-of-thought (CoT) data to fine-tune models and enhance their reasoning capabilities (Bismay et al., 2024; Tsai et al., 2024; Fang et al., 2025), mirroring trends in general LLM reasoning tasks. However, applying this paradigm to recommendation introduces fundamental challenges at both the tuning and inference stages. First, collecting high-quality CoT data for tuning is inherently difficult: (1) user feedback in recommendation is usually limited to final outcomes, with no access to the underlying reasoning process; and (2) the subjective, personalized nature of preferences (Tsai et al., 2024) makes it both costly and unreliable to manually annotate or synthesize CoT data aligned with individual users. Second, at inference time, these methods require generating explicit CoT reasoning, which is often verbose and incurs substantial latency—an important concern in real-world deployment. These limitations restrict the practical applicability of current methods. To address this, we pose a central question: Can we eliminate the need for explicit CoT reasoning at both the tuning and inference, while still unleashing the reasoning potential of LLM for recommendation?
A promising approach to achieve this is to shift from natural language reasoning to latent reasoning (Hao et al., 2024), where LLMs perform reasoning directly within their hidden representation space, thereby removing the need for explicit textual CoT reasoning. Moreover, hidden states offer much higher information density than textual tokens, allowing compact latent representations to encode complex reasoning processes, thereby reducing the need for lengthy reasoning chains (Su et al., 2025). This would enable more efficient inference. Despite their great potential, existing latent reasoning methods in general domains are not directly applicable to our goal, as they still rely on explicit CoT supervision to learn latent reasoning (Hao et al., 2024), such as by distilling CoT reasoning into latent reasoning.
This work explores learning latent reasoning without relying on any explicit CoT data, in an end-to-end optimization manner. Achieving this is challenging, as the only available signal for reasoning supervision is weak, coming solely from final user feedback, with no direct guidance on the reasoning process itself. Inspired by the recent success of reinforcement learning (RL) in learning explicit CoT reasoning strategies without CoT supervision, such as DeepSeek-R1-Zero (Guo et al., 2025), we investigate the use of RL to achieve this goal. However, training RL from scratch can be unstable and prone to collapse, particularly given the vast, high-dimensional space of latent reasoning. To mitigate this, we adopt a two-stage training strategy inspired by DeepSeek-R1 (Guo et al., 2025). In the first stage, we apply supervised fine-tuning to warm up the latent reasoning module, providing a strong initialization. In the second stage, we conduct pure RL training with a rule-based reward to promote exploration and further improve the reasoning ability.
Taking a step further, our RL approach builds on the GRPO algorithm (Shao et al., 2024), with task-specific modifications to the reward design. In its original main form, GRPO assigns a binary (0-or-1) reward by comparing the generated answer to ground truth, requiring full autoregressive answer generation for each sampled reasoning path—resulting in substantial computational overhead. To address this, we modify the reward to use the perplexity of the target item as a proxy, thereby eliminating the need for costly answer generation during training. This also produces a continuous reward signal, which could provide richer learning feedback and may facilitate more efficient optimization. Furthermore, we shift from a group-relative advantage to a batch-relative advantage by using the batch average reward as the baseline for advantage computation. This adjustment addresses the issue in the continuous reward setting, where the group-relative method could assign unreliable positive advantages, even if the entire group exhibits low-quality reasoning. Since our method operates with RL, we refer to it as Reinforced Latent Reasoning for Recommendation (LatentR3).
The main contribution of this work can be summarized as follows:
  • We highlight that latent reasoning provides a practical and latency-efficient method for integrating LLM reasoning into recommendation systems and propose to achieve latent reasoning in recommendations without relying on explicit CoT data.
  • We propose LatentR, a new method that enables latent reasoning with zero-shot CoT examples through a carefully designed RL framework. It features a novel reward formulation and an improved advantage computation mechanism to guide the optimization process of latent reasoning.
  • We conduct extensive experiments on real-world datasets, demonstrating that without CoT data, our latent reasoning approach can yield significant performance improvements.
2 Problem Definition
Let $\mathcal{D}$ denote the collected recommendation data, and let $(u,h,y)\in\mathcal{D}$ denote an instance, where $u$ is a user, $h$ is the user’s historical interactions, and $y$ is the next item the user interacts with. Both $h$ and $y$ are described using textual information (e.g., item titles). To leverage the capabilities of LLMs, we reformulate the recommendation task as a natural language problem. Specifically, for each data point $(u,h,y)$, we convert the historical interactions into a textual prompt $x$, which is then input to the LLM to generate a next-item recommendation. We can also represent a data point by $(x,y)$. Given that user preferences are implicitly and intricately embedded in the historical data, we avoid asking the LLM to produce the recommendation directly. Instead, we introduce a slow thinking process that encourages the model to first perform intermediate reasoning before generating the final output. This can be formalized as:
$x\xrightarrow{LLM(x)}r\xrightarrow{LLM(x,r)}\hat{y},$ (1)
where $r$ denotes the LLM’s intermediate reasoning output, i.e., thoughts, and $\hat{y}$ is the final predicted item. Importantly, we aim for this reasoning process to be efficient and learnable without requiring additional supervision in the form of explicit reasoning data ( CoT annotations). In other words, we train the LLM to perform recommendation-oriented reasoning directly from $\mathcal{D}$, without relying on externally annotated CoT data.
3 Methodology
Figure 1: An illustration of the model architecture and training strategy of the proposed LatentR3. On the architectural side, a LatentRATT layer is introduced to generate latent reasoning continuous tokens. The training follows a two-stage framework: the first stage performs warm-up tuning via supervised fine-tuning (SFT), while the second stage applies RL based on a modified GRPO to further optimize the reasoning.
We propose achieving inference-efficient, explicit CoT-free reasoning through reasoning in the latent space and leveraging reinforcement learning to enable reasoning learning. This forms our method, Reinforced Latent Reasoning for Recommendation (LatentR3). As shown in Figure 1, LatentR3 has key designs in both the reasoning architecture and the learning strategy side:
  • Latent Reasoning Architecture. To enable latent reasoning, LatentR3 adds an attention layer on top of the LLM’s final decoding layer, denoted by “LatentRATT". This layer extracts information from the final hidden states of the LLM to generate a latent reasoning token aligned with the LLM’s input embedding space, in an autoregressive manner. Together, the LLM and this generation layer constitute the full reasoning model.
  • Reinforced Learning. Given the lack of direct supervision for the reasoning process, the model must learn latent reasoning from final output signals. To enable the model to effectively learn latent reasoning, we use a reinforcement learning approach consisting of two stages: 1) Warm-up Fine-Tuning, which warms up the latent reasoning model through supervised fine-tuning, providing a meaningful initialization for the subsequent RL stage, and 2) RL-based Tuning, which applies pure RL to encourage exploration of reasoning and further enhance its reasoning capabilities. The RL stage is implemented using the GRPO method, with customized designs for sampling, reward, and advantage estimation to align with our specific requirements.
Next, we present the details of our latent reasoning architecture, followed by an explanation of the reinforced learning method.
3.1 Latent Reasoning Architecture
Before generating the final answer, we guide the LLM to first produce reasoning tokens in a continuous latent space to simulate slow thinking. Unlike prior work that directly uses hidden states from the LLM’s decoding layers as latent thoughts within the input embedding space, we introduce an additional attention layer on top of the final decoding layer to explicitly generate latent thought tokens. This layer serves two key purposes: (1) it functions as a dedicated reasoning token generator that aggregates contextual information to produce coherent latent thoughts, and (2) it ensures that the generated tokens are aligned with the LLM’s input embedding space.
Formally, given a data point converted into the prompt format $(x,y)$, at the $i$-th reasoning token generation, we first obtain the hidden states from the LLM’s final decoding layer by inputting $x$ along with all previously generated latent reasoning tokens. We then apply our additional attention layer to these hidden states and take the last-position output as the next latent reasoning token. Formally,
$H_{i-1}=LLM^{-1}(x,r_{1},...,r_{i-1}),$ (2)
$r_{i}=\text{LatentRATT}(H_{i-1})[-1],$ (3)
where $LLM^{-1}(\cdot)$ denotes the output of the final decoding layer of the LLM, $r_{1},\dots,r_{i-1}$ are the previously generated latent reasoning tokens, and $H_{i-1}$ is the corresponding sequence of hidden states produced by $LLM^{-1}(\cdot)$. “LatentRATT” refers to our additional attention layer for generating latent reasoning tokens. Finally, $r_{i}\in\mathbb{R}^{d}$ denotes the $i$-th generated latent reasoning token, where $d$ is the dimension size of the LLM’s text token embedding.
Final Generation. After generating $N$ latent reasoning tokens to form the final reasoning thought $r=[r_{1},\dots,r_{N}]$, we concatenate $x$ and $r$ as the input to the LLM for next-item prediction.
3.2 Reinforced Learning
Inspired by DeepSeek-R1, instead of performing RL from scratch, we first use supervised fine-tuning to warm up the model, providing a strong initial model for RL tuning. We then apply the pure RL to explore better reasoning, forming a two-stage tuning strategy. Next, we introduce each of these stages in detail.
3.2.1 Warm-up Fine-Tuning
Training RL from scratch can lead to instability and potential collapse, especially considering the vast, high-dimensional space of latent reasoning. To mitigate this, we first perform SFT to warm up the reasoning model, providing a solid initialization. Specifically, we fine-tune the full model using a standard next-token prediction task, enabling the model to generate meaningful latent reasoning tokens that enhance recommendation performance. The tuning objective is formulated as follows:
$\text{min}_{\Omega}\quad L_{warm}=-\sum_{(x,y)\in\mathcal{D}}\sum_{i=1}^{|y|}logP_{\theta}(y_{i}|x,r,y_{<i}),$ (4)
where $y_{i}$ denotes the $i$-th token of $y$, $|y|$ denotes the total number of tokens in $y$, and $P_{\theta}(y_{i}|x,r,y_{<i})=\text{LLM}(x,r,y_{<i})$ denotes the predicted probability of $y_{i}$ given the previous generations $y_{<i}$, latent reasoning $r$, and $x$; $\theta$ represents the total model parameters including those of the original LLM and our “LatentRATT” layer, and $L_{warm}$ denotes the total loss.
3.2.2 RL-based Tuning
Although the model may have gained some preliminary latent reasoning capabilities through SFT, its performance remains insufficient, as it might focus solely on data fitting rather than exploring optimal reasoning paths, leading to suboptimal outcomes. We consider leveraging RL training to facilitate exploration of more diverse reasoning paths, using the SFT results from the first stage as the initialization for further learning.
Given the demonstrated success of GRPO in DeepSeek-R1, we adopt it as the foundation for implementing our RL training. Generally, GRPO operates in four steps: (1) sampling a set of outputs from the old policy model; (2) computing a reward for each output; (3) calculating the advantage by comparing the group average reward; and (4) performing policy updates based on the advantage. To better align with our latent reasoning approach and enhance training efficiency, we introduce specific modifications to the first three steps. We now elaborate on these modifications:
$\bullet$ Sampling. Since our latent reasoning operates in a continuous space, it cannot be directly sampled like discrete textual tokens in standard GRPO. To address this, we leverage the Reparameterization Trick (Kingma et al., 2013, 2015), treating the generated latent reasoning vector as the mean of a Gaussian distribution. For each training sample $(x,y)$ with its corresponding latent reasoning vector $r$, we draw $K$ samples of reasoning. Formally, the $k$-th sampled reasoning ($k\neq 0$) is sampled as:
${r}^{k}=r+\epsilon,\quad\epsilon\sim\mathcal{N}(0,\sigma^{2}),$ (5)
where $\mathcal{N}(0,\sigma^{2})$ is a Gaussian distribution with zero mean and variance $\sigma^{2}$, $\epsilon$ denotes the sampled noise, and $\sigma$ can be thought of as a hyperparameter controlling the noise strength. Additionally, we include the original $r$ as the 00-th sample, i.e., $r^{0}=r$.
$\bullet$ Reward Design. After obtaining multiple latent reasoning samples in diverse directions, we evaluate which direction yields better outcomes by computing a reward for each. In the original GRPO, autoregressive decoding is used to generate the final output for each sample, and rewards are computed based on the quality of these outputs. However, this process is computationally expensive due to the need for full autoregressive decoding. To improve training efficiency, we instead propose using the model’s perplexity (PPL) on the ground-truth answer as a proxy for the reward. This approach allows for faster evaluation.
Specifically, we construct the LLM input by combining the original prompt, each sampled latent reasoning vector, and the ground-truth answer. We then compute the model’s perplexity for predicting the ground-truth tokens and use its negative value as the reward to evaluate each latent reasoning direction. Formally, the reward for the $k$-th sampled latent reasoning $r^{k}$ is defined as:
$s^{k}=-exp(-\frac{1}{|y|}\sum_{i=1}^{T}log\pi_{\theta}(y_{i}|x,r^{k},y_{<i})).$ (6)
Here, $\pi_{\theta}$ represents the current policy model (i.e., the current reasoning model), and $\pi_{\theta}(y_{i}|x,r^{k},y_{<i})$ denotes the predicted probability for the $i$-th token in $y$. The rewards corresponding to the group of sampled reasoning instances are denoted as $S=[s^{0},\dots,s^{K}]$. Notably, the rewards are continuous rather than discrete.
$\bullet$ Advantage Design. After obtaining the reward scores, we compute the advantage. In the original GRPO, the advantage for each sampled reasoning instance is computed by comparing its reward to the average reward within the same sampling group. However, since our rewards are continuous, this approach may still yield positive advantages even when all samples in the group are of low quality, thus producing unreliable advantage signals. To mitigate this issue, we instead compute the advantage by comparing each reward to the batch-level average reward. Specifically, for a sampled reasoning $r^{k}$ with reward score $s^{k}$, the advantage is computed as:
$\begin{split}A^{k}=\frac{s^{k}-\bar{s}_{batch}}{\|\mathbf{S}_{batch}-\bar{s}_{batch}\|},\end{split}$ (7)
[…]
$\bullet$ Policy Update. After computing the advantage, we apply the GRPO loss to update our model. The optimization objective is formulated as follows:
$L_{\text{GRPO}}=\sum_{(x,y)\in\mathcal{D}}-\frac{1}{K}\sum_{k=1}^{K}\frac{1}{|y|}\sum_{i=1}^{|y|}\frac{\pi_{\theta}(y_{i}\mid x,r^{k},y_{<i})}{\pi_{\text{old}}(y_{i}\mid x,r^{k},y_{<i})}\cdot A^{k}-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{ref}).$ (8)
Notably, to reduce computational cost, we update only the “LatentATT” layer while keeping the original LLM layers frozen. As a result, the second KL-divergence term, $\mathbb{D}_{KL}(\pi_{\theta}|\pi_{ref})$, becomes zero in the GRPO implementation.
4 Experiments
In this section, we conduct experiments to answer the following research questions: RQ1: How does the overall recommendation performance of our proposed LatentR3 model compare to existing recommendation methods? RQ2: Where do the improvements of LatentR3 come from? RQ3: How do different design choices and reasoning hyperparameters affect the effectiveness of our method?
4.1 Experimental Settings
Dataset.
In line with existing works, we conduct experiments on multiple domain-specific datasets from the Amazon review corpus, specifically focusing on three sub-datasets: Toys, CDs, and Games. Following the standard preprocessing procedure, we apply 5-core filtering to ensure data quality. Additionally, we adopt several temporal filtering strategies to maintain a manageable dataset size for RL training while preserving data quality—such as ensuring each dataset contains more than 5,000 items. More detailed preprocessing steps and dataset statistics are provided in Appendix A.
Compared Methods.
We consider the top-N recommendation task. To evaluate the effectiveness of our method, we compare it against both 1) representative traditional sequential recommendation models, including Caser (Tang and Wang, 2018), GRU4Rec (Hidasi and Karatzoglou, 2018), and SASRec (Kang and McAuley, 2018); and 2) LLM-based recommendation approaches, including AlphaRec (Sheng et al., 2025), BIGRec (Bao et al., 2025), and D3 Bao et al. (2024). AlphaRec is a state-of-the-art (SOTA) method that leverages LLM-generated embeddings to enhance the recommendation model, while D3 is an SOTA method that directly leverages LLMs as generative recommenders. More details about the baseline can be found in Append B. For our LatentR3, we implement it on both BIGRec and D3. Notably, existing LLM reasoning methods for recommendation cannot be selected for comparison because: (1) they are not designed for top-N recommendation tasks, and (2) their training relies on explicit CoT supervision, which is difficult to obtain in our setting. Similarly, we exclude general-domain latent reasoning methods for comparison, as they still depend on CoT data for learning. However, in our ablation study, we will include a variant without RL training, which can be viewed as an adaptation of the general latent reasoning method to our setting. The implementation details, including hyperparameter tuning, can be found in Appendix C.
Evaluation settings.
We evaluate the top-N recommendation effectiveness using Hit Ratio (HR@N) and Normalized Discounted Cumulative Gain (NDCG@N), with N set to 5 and 10. For brevity, we denote HR@5 and NDCG@5 as H@5 and N@5, respectively, in tables and figures; the same applies to the ‘@10’ cases.
4.2 Main Results (RQ1)
Table 1: Top-N recommendation performance comparison between our LatentR3 and baseline methods. “RI” denotes the relative improvement of our best-performing version (latentR3 applied to D3) over each baseline across all datasets and metrics. By default, the comparison is made against each column. For the BIGRec column specifically, the RI is calculated using LatentR3 applied to BIGRec. The best results are bold.
Dataset Methods Traditional LLM-based
Caser GRU4Rec SASRec AlphaRec BIGRec +LatentR3 D3 +LatentR3
Toys H@5 0.0251 0.0417 0.0601 0.0579 0.0701 0.0821 0.0830 0.0898
H@10 0.0384 0.0564 0.0760 0.0893 0.0931 0.1107 0.1026 0.1152
N@5 0.0170 0.0305 0.0458 0.0347 0.0508 0.0600 0.0610 0.0670
N@10 0.0214 0.0352 0.0510 0.0448 0.0582 0.0693 0.0674 0.0752
CDs H@5 0.0469 0.0481 0.0841 0.0479 0.0757 0.0934 0.1122 0.1137
H@10 0.0689 0.0669 0.1054 0.0774 0.0929 0.1160 0.1272 0.1327
N@5 0.0312 0.0365 0.0622 0.0278 0.0616 0.0754 0.0906 0.0915
N@10 0.0382 0.0425 0.0691 0.0373 0.0672 0.0826 0.0955 0.0977
Games H@5 0.0324 0.0322 0.0416 0.0558 0.0461 0.0580 0.0608 0.0716
H@10 0.0538 0.0517 0.0633 0.0893 0.0709 0.0870 0.0860 0.1006
N@5 0.0211 0.0207 0.0280 0.0397 0.0334 0.0413 0.0423 0.0507
N@10 0.0280 0.0270 0.0350 0.0515 0.0414 0.0506 0.0505 0.0601
RI 170.8% 121.0% 52.3% 77.5% 21.8% - 10.4% -
Table 1 presents the overall performance comparison between our method and the baselines, including both traditional (Caser, GRU4Rec, SASRec) and LLM-based (AlphaRec, BIGRec, D3) approaches. From the results, we draw the following conclusions:
  • The best version of our method (LatentR3 applied to D3) demonstrates superior performance compared to all existing approaches across all metrics on all three datasets, consistently validating the effectiveness of our proposed methodology.
  • Focusing on the comparison among LLM-based methods, those that use LLMs as recommenders (BIGRec and D3) outperform the method that leverages LLMs to enhance traditional models (AlphaRec) on the Toys and CDs datasets, but underperform on the Games dataset. After applying our LatentR3, the performance of both BIGRec and D3 improves significantly, with BIGRec achieving a relative improvement (RI) of 21.8% and D3 achieving an RI of 10.4%. Furthermore, in most cases, these enhanced models surpass AlphaRec. These results demonstrate that 1) incorporating latent reasoning substantially improves the performance of the LLM-based methods that leverage LLMs as the recommender; 2) our latent reasoning method can be applied to different existing LLM-based methods.
  • The best-performing LLM-based baselines generally outperform the strongest traditional baselines, consistent with findings in prior work. This advantage is largely attributed to the superior capabilities of LLMs, such as their ability to model rich world knowledge.
4.3 In-depth Analysis (RQ2 & RQ3)
In this section, we first analyze the performance improvements of our method across different item groups to understand where the gains originate. We then investigate the impact of various factors on the method’s effectiveness, starting with ablation studies to assess the contribution of each design component, followed by an analysis of the reasoning length. Lastly, we compare our modified GRPO with the original GRPO method. To reduce the computational cost (RL), we limit our analysis to the Toys and CDs datasets, except for the study analyzing the source of performance gains.
4.3.1 Performance over Popular and Unpopular Items
Figure 2: Performance improvement of LatentR3 over BIGRec on both popular and unpopular items.
For latent reasoning, it is difficult to directly demonstrate why it leads to performance improvements, as can be done with explicit CoT. Therefore, we consider providing indirect evidence. Intuitively, reasoning should be more beneficial for the more challenging aspects of recommendation, where it is expected to deliver greater gains. Long-tail (i.e., unpopular) items are typically more difficult to recommend accurately, while existing methods already perform well on popular items. To test this hypothesis, we examine whether our method yields greater improvements on unpopular items. Specifically, we compare the relative performance gains of our method over BIGRec across items with different popularity levels. The results are shown in Figure 2 (see Appendix D for more detailed results and experimental settings). As shown, our method achieves significantly larger improvements on unpopular items, suggesting that the incorporation of reasoning is particularly beneficial in more challenging recommendation scenarios.
Table 2: Ablation study results on the Toys and CDs datasets. The best results are highlighted in bold.
Method Toys CDs
H@5 H@10 N@5 N@10 H@5 H@10 N@5 N@10
LatentR3 0.0821 0.1107 0.0600 0.0693 0.0934 0.1160 0.0754 0.0826
w/o Reasoning 0.0701 0.0931 0.0508 0.0582 0.0757 0.0929 0.0616 0.0672
w/o LatentRATT 0.0772 0.1040 0.0574 0.0661 0.0705 0.0875 0.0568 0.0624
w/o RL (only SFT) 0.0804 0.1067 0.0584 0.0669 0.0830 0.1005 0.0662 0.0719
w/o Batch Advantage 0.0812 0.1083 0.0589 0.0676 0.0828 0.1002 0.0661 0.0718
4.3.2 Ablation Study
Our method features key innovations in both the model architecture and the learning algorithm. To assess the contribution of each component, we conduct ablation studies on both aspects. Specifically, we evaluate the following variants of our method: 1) w/o Reasoning, which removes the latent reasoning component entirely. 2) w/o LatentRATT, which omits the LatentRATT layer and directly uses the final hidden states from the LLM to construct the latent reasoning token; 3) w/o RL, which trains the reasoning model using only supervised fine-tuning (SFT), without RL; 4) w/o Batch Advantage, which replaces our batch-based advantage estimation with the original GRPO setting that computes advantages with a group average reward. We compare our default method with these variants on the Toys and CDs datasets. Notably, regarding the reward design, we would study it later when compared to the original GRPO.
The comparison results are summarized in Table 2. From the architectural perspective, removing the LatentRATT layer leads to a significant performance drop. On the CDs dataset, this variant even underperforms the "w/o Reasoning" version, underscoring the importance of LatentRATT in generating meaningful and space-aligned reasoning outputs. From the learning algorithm side, the "w/o RL" variant—which only uses SFT—performs better than the "w/o Reasoning" version, indicating that SFT can partially enable latent reasoning. However, its performance remains inferior to the full method with RL, confirming that effective latent reasoning learning relies on reinforcement training. Furthermore, replacing the batch-based advantage with the original GRPO’s advantage (as in "w/o Batch Advantage") results in comparable or worse performance than the "w/o RL" version, demonstrating the necessity of batch-based reward estimation for stable RL learning under our reward design. Overall, these findings highlight the critical roles of both our architectural and learning algorithm designs in the method’s effectiveness.
4.3.3 Influence of Reasoning Length
Figure 3: Impact of Reasoning Length on LatentR3 Performance.
[…]
4.3.4 Comparison with the Original GRPO
We also compare our modified GRPO with the original GRPO, which relies on generating complete answers for reward computation for group sampling, to evaluate its effectiveness. To improve training efficiency, we propose using the model’s perplexity (PPL) on the ground-truth answer as a proxy for the reward. This yields a continuous reward signal and motivates adjustments to the advantage computation. To assess the impact of these modifications, we implement a variant of our method using the original GRPO to implement the RL. The comparison results on the CDs dataset are shown in Figure 4, with additional results and implementation details provided in Appendix F. In terms of performance, our method achieves results comparable to the variant using original GRPO—slightly outperforming it on the CDs dataset while performing slightly worse on the Toys dataset (see Appendix F). In terms of training efficiency, our method is significantly faster, reducing the training cost to approximately 1/5 of that of the original GRPO-based method. These findings validate the rationale and effectiveness of our proposed modifications.
Figure 4: Comparison to the original GRPO regarding Performance (Left) and efficiency (Right).
5 Related Work
Our work relates to LLM-based recommendation and latent reasoning. Due to space constraints, we briefly discuss LLM-based recommendation and its associated reasoning methods here, with a more comprehensive review available in Appendix G. The recent widespread success of LLMs has sparked growing interest in their application to recommendation, giving rise to three main LLM-based recommendation paradigms: (1) in-context learning (Gao et al., 2023; Sun et al., 2024; Wang et al., 2021), (2) agent-based frameworks (Wang et al., 2023; Zhang et al., 2024a, b), and (3) fine-tuning-based methods (Bao et al., 2025; Zhang et al., 2024c). Among these, the fine-tuning-based solution has received the most attention due to its ability to more effectively align LLMs with the task, and our work falls into this category. In this direction, only a few studies have explored reasoning, with existing works primarily aiming to enhance explicit chain-of-thought (CoT) reasoning via fine-tuning, such as RecSAVER (Tsai et al., 2024), ReasoningRec (Bismay et al., 2024), Reasoning4Rec (Fang et al., 2025), and Exp3rt (Kim et al., ). However, these approaches rely heavily on CoT data, which is costly to obtain and often difficult to ensure in quality. Moreover, the use of explicit CoT reasoning typically incurs higher inference latency. In contrast, our work introduces latent reasoning for LLM-based recommendation, which eliminates the need for CoT supervision and enables more efficient inference. As for latent reasoning, only two concurrent works (Tang et al., 2025; Zhang et al., 2025) have explored it, but neither focuses on LLM-based recommendation, and our approach to latent reasoning is fundamentally different and specifically tailored for LLM-based recommendation.
6 Conclusion
In this work, we proposed LatentR3, a novel LLM-based recommendation framework that replaces explicit chain-of-thought reasoning with compact latent reasoning representations. By integrating architectural innovations such as the LatentRATT layer with a two-stage reinforcement learning strategy, LatentR3 enables efficient and effective preference reasoning without relying on costly CoT supervision. Extensive experiments show that our method significantly enhances the performance of existing LLM-based recommendation models, while introducing only minimal additional inference cost—equivalent to generating a single extra token. These results highlight the potential of latent reasoning as a practical and scalable alternative for LLM-based recommendation.
Limitations and Future Works.
Due to current computational resource constraints, our experiments are limited to relatively small datasets. In the future, with access to greater computational resources, we plan to explore the effectiveness of our method on larger-scale datasets. Additionally, our method currently faces challenges when generating long latent reasoning sequences. Future work will focus on developing more effective and efficient approaches to learn latent reasoning with extended reasoning lengths. Last but not least, compared to explicit CoT, latent reasoning offers less interpretability. We intend to investigate ways to improve the explainability of latent reasoning, which could be crucial for its broader adoption and practical use.
Appendix A Dataset Pre-processing
Our experiments were conducted on the Amazon review dataset, with primary evaluations performed on three domain-specific subsets: Toys, CDs, and Games. Following established methodology, we preprocessed each subset to achieve 5-core data integrity. To address the computational demands of LLM training while maintaining dataset validity, we implemented a dynamic temporal partitioning strategy: We iteratively process the dataset within a sliding time window, starting from October 2017 to October 2018. After applying 5-core filtering, if the resulting item count falls below 5,000, the start time is shifted back by 3 months (e.g., to July 2017). This adjustment repeats until the processed dataset contains more than 5,000 items, with the end time fixed at October 2018 throughout the process. Key dataset statistics are summarized in the accompanying table3. To align with real-world scenarios, we adopt a temporal split of the processed datasets into training, validation, and testing sets based on timestamps, following an 8:1:1 ratio. In alignment with comparative baseline requirements across all experimental conditions, we standardized the maximum sequence length at 10 for consistency in temporal pattern analysis.
Table 3: Dataset statistics.
Dataset Train Valid Test Item
Toys 53898 6737 6738 6299
CDs 49251 6156 6158 5841
Games 75175 9397 9397 5308
Appendix B Compared Methods
We consider the top-N recommendation task. To evaluate the effectiveness of our method, we compare it against both representative traditional sequential recommendation models and state-of-the-art (SOTA) LLM-based recommendation approaches:
    Caser (Tang and Wang, 2018): This is a well-known sequential recommendation approach that utilizes Convolutional Neural Networks (CNNs) to capture sequential patterns and model user preferences.
    GRU4Rec (Hidasi and Karatzoglou, 2018): This is another widely recognized method that leverages Gated Recurrent Units (GRUs) to encode sequential patterns and model user preferences.
    SASRec (Kang and McAuley, 2018): This is a highly representative sequential recommendation method that utilizes a self-attention network to model user preferences, with an architecture resembling that of the decoder-only Transformer model.
    AlphaRec (Sheng et al., 2025): This is an SOTA LLM-based method that leverages LLM-generated embeddings to enhance the recommendation model.
    BIGRec (Bao et al., 2025): This is a representative LLM-based generative recommendation method that fine-tunes LLMs to generate next-item predictions, with specific designs to support full ranking.
    D3 Bao et al. (2024): This is a SOTA LLM-based generative recommendation method. It follows a similar fine-tuning process to BIGRec but differs in its inference strategy. Specifically, it introduces debiasing techniques during inference to enhance the quality of generated recommendations. Additionally, it includes an ensemble design with traditional models; however, we omit this component in our implementation to ensure a fair comparison.
Our method, LatentR3, is compatible with various LLM-based generative recommendation approaches. Accordingly, we implement it on both BIGRec and D3. Notably, existing LLM reasoning methods for recommendation cannot be selected for comparison because: (1) they are not designed for top-N recommendation tasks, and (2) their training relies on explicit CoT supervision, which is difficult to obtain in our setting. Similarly, we exclude general-domain latent reasoning methods for comparison, as they still depend on CoT data for learning. However, in our ablation study, we will include a variant of our method without RL training, which can be viewed as an adaptation of the general latent reasoning method to our setting.
Appendix C Implementation Details
For traditional recommendation models, we use the Adam optimizer and perform grid searches over learning rates 1e-2, 1e-3, 1e-4 and weight decay values 1e-4, 1e-5, 1e-6, and all models are trained using Binary Cross-Entropy loss with randomly sampled negative items. For LLM-based methods, we use Qwen2.5-1.5B Team (2024) as the backbone LLM. Supervised fine-tuning (SFT) is conducted using the AdamW optimizer, with learning rates selected from {3e-3, 3e-4, 3e-5}, and early stopping is applied with a patience of 1. During the reinforcement learning stage, we search learning rates within {1e-5, 1e-4, 5e-4}. All experiments are run on 2 NVIDIA A100 GPUs.
Appendix D Performance on Popular and Unpopular Items
Based on the frequency of items in the training set, we consider the top 20% of items as popular items and the other items as unpopular items. We then calculate the performance of BIGRec and LatentR3 for these two groups of items, and calculate the relative performance improvement in each group. The specific results are shown in Figure 5.
Figure 5: Performance improvement of LatentR3 over BIGRec on both popular and unpopular items.
Appendix E Influence of Reasoning Length
Figure 6: Impact of Reasoning Length on LatentR3 Performance
[…]
Appendix F Comparison with the Original GRPO
Figure 7: Comparison to the original GRPO regarding Performance (Left) and efficiency (Right).
To implement the variant with the original GRPO, after sampling the latent reasoning, we leverage the LLM to generate the final answer, and then exactly match the answer with the ground-truth to compute 0-1 reward. Lastly we leverage the group average for advantage computation. The experiments are conducted on both the Toys and CDs datasets. The results are summarized in Figure 7. In terms of performance, our method achieves results comparable to the variant using original GRPO—slightly outperforming it on the CDs dataset while performing slightly worse on the Toys dataset. In terms of training efficiency, our method is significantly faster, reducing the training cost to approximately 1/5 of that of the original GRPO-based method. These findings validate the rationale and effectiveness of our proposed modifications. Notably, the reported results for the original GRPO are based on full-model tuning during RL, as this setting is necessary to achieve the reported performance. Tuning only the LatentRATT layer, as done in the default LatentR3—results in significantly worse outcomes for GRPO. However, we observe that limiting GRPO to tuning the LatentRATT layer would only slightly reduce training cost (by approximately 50 minutes), indicating that our method remains substantially more efficient.
Appendix G Full Related Work
LLM-based Recommendation.
The recent widespread success of LLMs has sparked growing interest in their application to recommendation systems, giving rise to three main paradigms: (1) in-context learning (Gao et al., 2023; Sun et al., 2024; Wang et al., 2021), (2) agent-based frameworks (Wang et al., 2023; Zhang et al., 2024a, b), and (3) fine-tuning-based methods (Bao et al., 2025; Zhang et al., 2024c). Among these, the fine-tuning-based solution has received the most attention due to its ability to more effectively align LLMs with recommendation objectives, and our work falls into this category. Although significant efforts have been devoted to this direction, most existing methods overlook the role of reasoning in recommendation—particularly latent reasoning, which remains unexplored. Among the limited studies that investigate reasoning, existing works primarily focus on enhancing explicit chain-of-thought (CoT) reasoning through fine-tuning. RecSAVER (Tsai et al., 2024) and ReasoningRec (Bismay et al., 2024) leverage larger LLMs to generate CoT reasoning data for training smaller LLMs, aiming to improve their reasoning capabilities. Reasoning4Rec (Fang et al., 2025) decomposes the reasoning process into multiple steps, using user reviews as proxy supervision. Exp3rt (Kim et al., ) focuses on distilling the reasoning behavior of larger LLMs into smaller models. However, these approaches rely heavily on CoT data, which is costly to obtain and often difficult to ensure in quality. Moreover, the use of explicit CoT reasoning typically incurs higher inference latency. In contrast, our work introduces latent reasoning for LLM-based recommendation, which eliminates the need for CoT supervision and enables more efficient inference. As for latent reasoning, only two concurrent works (Tang et al., 2025; Zhang et al., 2025) have explored it, but neither focuses on LLM-based recommendation, and our approach to latent reasoning is fundamentally different and specifically tailored for LLM-based recommendation.
Latent reasoning in LLM.
Latent reasoning has been proposed to address two major limitations of explicit CoT reasoning: (1) high inference latency and (2) the excessive generation of non-essential tokens. Several efforts have been made to explore latent reasoning mechanisms. For example, Deng et al. (2023) introduced the implicit CoT by encoding reasoning directly into the model’s internal hidden states, while Hao et al. (2024) proposed treating these hidden states as thoughts in the input space to represent reasoning. However, both methods rely on supervision from explicit CoT data. Goyal et al. (2024) proposed using learnable <pause> tokens to represent reasoning steps, but this approach may suffer from low expressivity (Hao et al., 2024). Among the existing works, the method proposed in (Geiping et al., 2025) is most related to ours. It introduces a recurrent architecture for generating deep latent reasoning, but it is not designed using reinforcement learning and requires pertaining a LLM model with new architecture, making it incompatible with direct integration into existing LLMs.